OpenStudy (anonymous):

Perform the indicated operation g(a) = 3a + 1 f(a) = a3 + 2 Find g(a) + f(a)

OpenStudy (anonymous):

just add the two functions. g(a)+f(a)=3a+1+3a+2=6a+3

OpenStudy (anonymous):

HOW

OpenStudy (anonymous):

You can substitute g(a) for 3a+1 and f(a)for 3a+2. So the sum of g(a)+f(a) is equivalent to the sum of their substitutions. its like x=6 and y=7, x+y=6+7=13